# On the arithmetic of polynomial ideals

### v1

@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
Fix $n\geq 3$ and let $r\in\llbracket 3,n\rrbracket$. Assume as a contradiction that $\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$. By Lemma 4.12, $X^{\alpha},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta}\in G(\mathfrak{b})$ with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})=a_{I}$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})=a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha=a_{(\{1\}\cup\llbracket 3,r\rrbracket)\setminus I}$, where $I$ is a nonempty proper subset of $\{1\}\cup\llbracket 3,r\rrbracket$}. By symmetry, we can also assume without loss of generality that $\alpha\leq\beta$. A rough estimate ensures then that $\alpha<a_{r}$. In fact, if $\alpha\geq a_{r}$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta\geq 2\alpha\geq 2a_{r}$, which is impossible, as $a_{r}>a_{\llbracket 1,r-1\rrbracket}$. Thus, we can assume without loss of generality that, if $r\geq 4$, then $\alpha=a_{I}$ with $I\subseteq\{1\}\cup\llbracket 3,r-1\rrbracket$, while, if $r=3$, then $\alpha=a_{1}$. As a consequence, $\beta\geq a_{r}$.
It is also easy to observe that we can exclude that $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$ and $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$, i.e., that
@@
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}}\rangle.$ (12)
In fact, the generators of $\tilde{\mathfrak{b}}_{r}$ in equation (12) are of the form $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}$, with $J=\emptyset,\{1\}\cup\llbracket 3,r\rrbracket,I,(\{1\}\cup\llbracket 3,r\rrbracket)\setminus I$. Therefore, by the observation following equation (8), none of them divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ and this leads to the contradiction $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\notin\tilde{\mathfrak{b}}_{r}$.
@@
Now we distinguish two cases depending on whether $\alpha$ is equal or not to $a_{1}$, and for each case we consider three subcases.
CASE 1: $\alpha=a_{1}$ and $\beta=a_{\llbracket 3,r\rrbracket}$.
Subcase I: $\mathfrak{a}=\langle X^{a_{1}},Y^{a_{1}}\rangle$ and $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{l+a_{1}}Y^{m},X^{l}Y^{m+a_{1}},\dots\rangle,$ (13)
with $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$ and $1\leq l,m<a_{\llbracket 3,r\rrbracket}$.
We know from (8) that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, hence by Lemma 4.1 at least one of the generators in (13) must divide it. From remarks following (8), it can only happen that $X^{l+a_{1}}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or that $X^{l}Y^{m+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, for some $l,m$ satisfying the requirements above.
Say that there are $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$, $1\leq l,m<a_{\llbracket 3,r\rrbracket}$ and
$X^{l+a_{1}}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1},$ (14)
$m\leq a_{3}-a_{2}.$ (15)
@@
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l}Y^{m+a_{1}}$ and this leads to the contradiction $X^{l}Y^{m+a_{1}}\notin\tilde{\mathfrak{b}}_{r}$. To prove the last assumption we first notice that, being $l<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+a_{1}}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l}Y^{m+a_{1}},$ (16)
using (14), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1}$, implying that $a_{J}\geq 2a_{1}+a_{2}$ and so that $a_{J}\geq a_{3}$ by (9). But (16) also implies $a_{J}\leq m+a_{1}$ and, by (15), we reach the contradiction $a_{J}\leq a_{3}-a_{2}+a_{1}<a_{3}$ (recall that $a_{2}>a_{1}$).
On the other hand, assume that there are $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$, $1\leq l,m<a_{\llbracket 3,r\rrbracket}$ and
$X^{l}Y^{m+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (17)
$m\leq a_{3}-a_{2}-a_{1}.$ (18)
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l+a_{1}}Y^{m}$ and this leads to the contradiction $X^{l+a_{1}}Y^{m}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last assumption we first notice that, being $m<a_{3}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l+a_{1}}Y^{m}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l+a_{1}}Y^{m},$ (19)
@@
using (17), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l+a_{1}\leq a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{2}$, implying that $a_{J}\geq a_{2}$ and so that $a_{J}\geq a_{3}$ by (9). But (19) also implies $a_{J}\leq m$ and, by (18), we reach the contradiction $a_{J}\leq a_{3}-a_{2}-a_{1}<a_{3}$.
Subcase II: $X^{a_{1}},Y^{a_{1}},X^{i}Y^{j}\in G(\mathfrak{a})$ and $\mathfrak{b}=\langle X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}}\rangle$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},X^{i}Y^{j+a_{\llbracket 3,r\rrbracket}},\dots\rangle,$ (20)
with $i,j\in\mathbb{N}$ such that $i+j\geq a_{1}$ and $1\leq i,j<a_{1}$.
This case never occurs. In fact, for any choice of $i,j\in\mathbb{N}$ such that $i+j\geq{a_{1}}$ and $1\leq i,j<a_{1}$, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$ and this leads to the contradiction $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}\notin\tilde{\mathfrak{b}}_{r}$. To show that this last assumption is true we first notice that, being $j<a_{1}<a_{3}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},$ (21)
we get $a_{J}\leq j<a_{1}$, implying that $a_{J}=0$ (this is a trivial consequence of the definition of $J$), whence $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}>i+a_{\llbracket 3,r\rrbracket}$, a contradiction.
Subcase III: $X^{a_{1}},Y^{a_{1}},X^{i}Y^{j}\in G(\mathfrak{a})$ and $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{a_{1}}Y^{a_{\llbracket 3,r\rrbracket}},X^{a_{\llbracket 3,r\rrbracket}}Y^{a_{1}},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j},X^{i}Y^{j+a_{\llbracket 3,r\rrbracket}},X^{l+a_{1}}Y^{m},X^{l}Y^{m+a_{1}},\dots\rangle,$ (22)
with $i,j\in\mathbb{N}$ such that $i+j\geq a_{1}$ and $1\leq i,j<a_{1}$ and $l,m\in\mathbb{N}$ such that $l+m\geq a_{\llbracket 3,r\rrbracket}$ and $1\leq l,m<a_{\llbracket 3,r\rrbracket}$.
Arguing as in the previous subcase, also this case can never happen.
CASE 2: $\alpha\neq a_{1}$, i.e., $\alpha\geq a_{3}$ by Lemma 4.12(11). Under our assumption that $\alpha<a_{r}$ (and so $\beta\geq a_{r}$), this case is only possible if $r\geq 4$ (so also $n\geq 4$). As a consequence, we also have $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, where $\llbracket 4,r\rrbracket$ is well defined.
Subcase I: $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$ and $X^{\beta},Y^{\beta},X^{l}Y^{m}\in G(\mathfrak{b})$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{l+\alpha}Y^{m},X^{l}Y^{m+\alpha},\dots\rangle,$ (23)
with $l,m\in\mathbb{N}$ such that $l+m\geq\beta$ and $1\leq l,m<\beta$.
We know from (8) that $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, therefore by Lemma 4.1 at least one of the generators in (23) must divide it. From remarks following (8), it can only happen that $X^{l+\alpha}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or that $X^{l}Y^{m+\alpha}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, for some $l,m$ satisfying the requirements above. The latter case can be also discarded: being $\alpha\geq a_{3}$ and $m\geq 1$, $m+\alpha>a_{3}>a_{3}-a_{2}$.
We then have just one case to consider. Say that there are $l,m\in\mathbb{N}$ such that $l+m\geq\beta$, $1\leq l,m<\beta$ and
$X^{l+\alpha}Y^{m}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Therefore,
$l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha,$ (24)
$m\leq a_{3}-a_{2}.$ (25)
@@
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{l}Y^{m+\alpha}$ and this leads to the contradiction $X^{l}Y^{m+\alpha}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last fact we first notice that, being $l<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{l}Y^{m+\alpha}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{l}Y^{m+\alpha},$ (26)
@@
using (24), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq l\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha$, implying that $a_{J}\geq a_{1}+a_{2}+\alpha$. Recalling that $\alpha=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and that $r\geq 4$, by Lemma 4.10 we have then two possibilities: $a_{J}=a_{3}-a_{1}+\alpha$ or $a_{J}\geq a_{3}+\alpha$. But (26) and (25) also imply $a_{J}\leq m+\alpha\leq a_{3}-a_{2}+\alpha$ and in both the cases above we reach a contradiction.
Subcase II: $X^{\alpha},Y^{\alpha},X^{i}Y^{j}\in G(\mathfrak{a})$ and $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$, i.e.,
$\tilde{\mathfrak{b}}_{r}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{i+\beta}Y^{j},X^{i}Y^{j+\beta},\dots\rangle,$ (27)
with $i,j\in\mathbb{N}$ such that $i+j\geq\alpha$ and $1\leq i,j<\alpha$.
@@
Arguing as in the previous case, there must exist $i,j\in\mathbb{N}$ such that $i+j\geq\alpha$, $1\leq i,j<\alpha$ and
$X^{i+\beta}Y^{j}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}.$
Note in fact that, since $j+\beta>a_{r}>a_{3}>a_{3}-a_{2}$, $X^{i}Y^{j+\beta}\nmid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Then,
@@
But then, none of the generators of $\tilde{\mathfrak{b}}_{r}$ in (10) divides $X^{i}Y^{j+\beta}$ and this leads to the contradiction $X^{i}Y^{j+\beta}\notin\tilde{\mathfrak{b}}_{r}$. To prove this last assumption, we first notice that, being $i<a_{\llbracket 3,r\rrbracket}-a_{2}$, $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i}Y^{j+\beta}$. Moreover, if for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$,
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i}Y^{j+\beta},$ (30)

### v2

@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
Fix $n\geq 3$ and let $r\in\llbracket 3,n\rrbracket$. Assume as a contradiction that $\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$. By Corollary 4.3, we can assume $G(\mathfrak{a}),G(\mathfrak{b})\subseteq K[X,Y]$ without loss of generality. By Lemma 4.13, $X^{\alpha},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta}\in G(\mathfrak{b})$ with $\alpha=\operatorname{\mathsf{m-deg}}(\mathfrak{a})=a_{I}$ and $\beta=\operatorname{\mathsf{m-deg}}(\mathfrak{b})=a_{\{1\}\cup\llbracket 3,r\rrbracket}-\alpha=a_{(\{1\}\cup\llbracket 3,r\rrbracket)\setminus I}$, where $I$ is a nonempty proper subset of $\{1\}\cup\llbracket 3,r\rrbracket$}. By symmetry, we can also assume without loss of generality that $\alpha\leq\beta$. A rough estimate then ensures that $\alpha<a_{r}$. In fact, if $\alpha\geq a_{r}$, then $a_{\{1\}\cup\llbracket 3,r\rrbracket}=\alpha+\beta\geq 2\alpha\geq 2a_{r}$, which is impossible, as $a_{r}>a_{\llbracket 1,r-1\rrbracket}$. Thus, we can assume without loss of generality that, if $r\geq 4$, then $\alpha=a_{I}$ with $I\subseteq\{1\}\cup\llbracket 3,r-1\rrbracket$, while, if $r=3$, then $\alpha=a_{1}$. As a consequence, $\beta\geq a_{r}$.
CASE 0: $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$, $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$
In this case, we have
@@
$\tilde{\mathfrak{b}}_{r}=\mathfrak{a}\cdot\mathfrak{b}=\langle X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}},X^{\alpha}Y^{\beta},X^{\beta}Y^{\alpha},Y^{a_{\{1\}\cup\llbracket 3,r\rrbracket}}\rangle\text{.}$ (12)
Since $\alpha=a_{I}$ for some $I\subseteq{\{1\}\cup\llbracket 3,r\rrbracket}$ and $\beta=a_{{\{1\}\cup\llbracket 3,r\rrbracket}\setminus{I}}$, the remarks following equation (8) imply that none of the generators in (12) divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, against Lemma 4.1.
@@
Now we distinguish two further cases depending on whether $\alpha$ is equal or not to $a_{1}$, and for each case we consider several subcases according to the structure of $\mathfrak{a}$ and $\mathfrak{b}$.
CASE 1: $\alpha=a_{1}$ and $\beta=a_{\llbracket 3,r\rrbracket}$. In particular, $X^{a_{1}},Y^{a_{1}}\in G(\mathfrak{a})$ and $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}}\in G(\mathfrak{b})$.
Subcase I: $\mathfrak{a}=\langle X^{a_{1}},Y^{a_{1}}\rangle$
Since $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, by Lemma 4.1 at least one of the minimal generators of $\mathfrak{a}\cdot\mathfrak{b}$ must divide it. By CASE 0, there must exist $k,\ell\in\mathbb{N}$ such that $X^{k}Y^{\ell}\in G(\mathfrak{b})$ and either $X^{k+a_{1}}Y^{\ell}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or $X^{k}Y^{\ell+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Note that $k+\ell\geq\operatorname{\mathsf{m-deg}}(\mathfrak{b})=a_{\llbracket 3,r\rrbracket}$ and, since $X^{a_{\llbracket 3,r\rrbracket}},Y^{a_{\llbracket 3,r\rrbracket}}\in G(\mathfrak{b})$, necessarily $1\leq k,\ell<a_{\llbracket 3,r\rrbracket}$.
Assume that $X^{k+a_{1}}Y^{\ell}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Therefore,
$k$ $\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1},$ (13)
$\ell$ $\leq a_{3}-a_{2}.$ (14)
@@
We claim that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{k}Y^{\ell+a_{1}}$, implying $X^{k}Y^{\ell+a_{1}}\in{(\mathfrak{a}\cdot\mathfrak{b})\setminus\tilde{\mathfrak{b}}_{r}}$, a contradiction. To prove our claim, we first notice that, since $k<a_{\llbracket 3,r\rrbracket}-a_{2}$, we have $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{k}Y^{\ell+a_{1}}$. Moreover, if
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{k}Y^{\ell+a_{1}}\text{\ for some\ }J\subseteq\{1\}\cup\llbracket 3,r\rrbracket,$ (15)
@@
then, using (13), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq k\leq a_{\llbracket 3,r\rrbracket}-a_{2}-a_{1}$, implying $a_{J}\geq 2a_{1}+a_{2}$ and so $a_{J}\geq a_{3}$, by (9). But (15) also implies $a_{J}\leq\ell+a_{1}$ and, by (14), we reach the contradiction $a_{J}\leq a_{3}-a_{2}+a_{1}<a_{3}$.
On the other hand, assume that $X^{k}Y^{\ell+a_{1}}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Therefore,
$k$ $\leq a_{\llbracket 3,r\rrbracket}-a_{2},$ (16)
$\ell$ $\leq a_{3}-a_{2}-a_{1}.$ (17)
@@
We claim that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{k+a_{1}}Y^{\ell}$, leading to the contradiction $X^{k+a_{1}}Y^{\ell}\notin\tilde{\mathfrak{b}}_{r}$. To prove this claim, we first notice that, since $\ell<a_{3}-a_{2}$, we have $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{k+a_{1}}Y^{\ell}$. Moreover, if
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{k+a_{1}}Y^{\ell}\text{\ for some\ }J\subseteq\{1\}\cup\llbracket 3,r\rrbracket,$ (18)
then, using (16), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq k+a_{1}\leq a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{2}$, implying $a_{J}\geq a_{2}$ and so $a_{J}\geq a_{3}$, by (9). But (18) also implies $a_{J}\leq\ell$ and, by (17), we reach the contradiction $a_{J}\leq a_{3}-a_{2}-a_{1}<a_{3}$.
Subcase II: $\mathfrak{a}\supsetneq\langle X^{a_{1}},Y^{a_{1}}\rangle$
As $X^{a_{1}},Y^{a_{1}}\in G(\mathfrak{a})$, there must exist $i,j\in\mathbb{N}$ such that $i+j\geq{a_{1}}$, $1\leq i,j<a_{1}$ and $X^{i}Y^{j}\in G(\mathfrak{a})$. We claim that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$, leading to the contradiction $X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}\in{(\mathfrak{a}\cdot\mathfrak{b})\setminus\tilde{\mathfrak{b}}_{r}}$. To prove our claim, we first notice that, since $j<a_{1}<a_{3}-a_{2}$, we have $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$. Moreover, if $X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i+a_{\llbracket 3,r\rrbracket}}Y^{j}$ for some $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$, then we get $a_{J}\leq j<a_{1}$, implying $a_{J}=0$, whence $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}=a_{\{1\}\cup\llbracket 3,r\rrbracket}>i+a_{\llbracket 3,r\rrbracket}$, a contradiction.
CASE 2: $\alpha\neq a_{1}$, i.e., $\alpha\geq a_{3}$ by Lemma 4.13(11). As we already observed in the initial part of the proof, under our assumptions this can happen only if $r\geq 4$ (in particular, $n\geq 4$). As a consequence, we also have $\beta\leq a_{\{1\}\cup\llbracket 4,r\rrbracket}$, where $\llbracket 4,r\rrbracket$ is well defined. Recall that $X^{\alpha},Y^{\alpha}\in G(\mathfrak{a})$ and $X^{\beta},Y^{\beta}\in G(\mathfrak{b})$.
Subcase I: $\mathfrak{a}=\langle X^{\alpha},Y^{\alpha}\rangle$
As $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\in\tilde{\mathfrak{b}}_{r}$, one of the minimal generators of $\mathfrak{a}\cdot\mathfrak{b}$ must divide it. By CASE 0, there must exist $X^{k}Y^{\ell}\in G(\mathfrak{b})$ (in particular, $k+\ell\geq\beta$ and $1\leq k,\ell<\beta$) such that either $X^{k+\alpha}Y^{\ell}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$ or $X^{k}Y^{\ell+\alpha}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. We can immediately discard the latter case: since $\alpha\geq a_{3}$ and $\ell\geq 1$, we always have $\ell+\alpha>a_{3}>a_{3}-a_{2}$.
Therefore, we must have $X^{k+\alpha}Y^{\ell}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, i.e.,
$k$ $\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha,$ (19)
$\ell$ $\leq a_{3}-a_{2}.$ (20)
We claim that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{k}Y^{\ell+\alpha}$, leading to the contradiction $X^{k}Y^{\ell+\alpha}\in(\mathfrak{a}\cdot\mathfrak{b})\setminus\tilde{\mathfrak{b}}_{r}$. To prove this claim we first notice that, since $k<a_{\llbracket 3,r\rrbracket}-a_{2}$, we have $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{k}Y^{\ell+\alpha}$. Moreover, if
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{k}Y^{\ell+\alpha}\text{\ for some\ }J\subseteq\{1\}\cup\llbracket 3,r\rrbracket,$ (21)
@@
then, using (19), we get $a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}\leq k\leq a_{\llbracket 3,r\rrbracket}-a_{2}-\alpha$, implying that $a_{J}\geq a_{1}+a_{2}+\alpha$. Recalling that $\alpha=a_{I}$ for some $I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ and that $r\geq 4$, by Lemma 4.11 we have two possibilities: either $a_{J}=a_{3}-a_{1}+\alpha$ or $a_{J}\geq a_{3}+\alpha$. But (21) and (20) also imply $a_{J}\leq\ell+\alpha\leq a_{3}-a_{2}+\alpha$ and in both the above cases we reach a contradiction.
Subcase II: $\mathfrak{b}=\langle X^{\beta},Y^{\beta}\rangle$
@@
Arguing as in the previous subcase, there must exist $X^{i}Y^{j}\in G(\mathfrak{a})$ such that $i+j\geq\alpha$, $1\leq i,j<\alpha$ and $X^{i+\beta}Y^{j}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Indeed, since $j+\beta>a_{r}>a_{3}>a_{3}-a_{2}$, we can discard the case $X^{i}Y^{j+\beta}\mid X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$. Thus,
@@
We claim that none of the minimal generators of $\tilde{\mathfrak{b}}_{r}$ divides $X^{i}Y^{j+\beta}$, leading to the contradiction $X^{i}Y^{j+\beta}\in(\mathfrak{a}\cdot\mathfrak{b})\setminus\tilde{\mathfrak{b}}_{r}$. To prove our claim, we first notice that, since $i<a_{\llbracket 3,r\rrbracket}-a_{2}$, we have $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\nmid X^{i}Y^{j+\beta}$. Moreover, if
$X^{a_{\{1\}\cup\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\mid X^{i}Y^{j+\beta}\text{\ for some\ }J\subseteq\{1\}\cup\llbracket 3,r\rrbracket,$ (24)